OpenStudy (anonymous):

F(x) = sqrt(x^2+2x) x>0 Find the inverse of the function.

OpenStudy (anonymous):

Swap x and F(x) then solve for F(x)

OpenStudy (anonymous):

I tried that but I didn't know where to go from there, I ended up with x = sqrt(y^2+2y) then x^2 = y^2 +2y, then x^2-y^2 = 2y and now I am at (x+y)(x-y) = 2y, I dont know where to go from here

OpenStudy (anonymous):

to get y by itself

OpenStudy (anonymous):

complete the square... y^2 + 2y + ?

OpenStudy (anonymous):

y^2 +2y +1 = (y+1)^2

OpenStudy (anonymous):

Where did the 1 come from though?

OpenStudy (anonymous):

so if we add 1 to both sides we get... x^2 +1 = y^2 +2y +1 x^2 +1 = (y+1)^2

OpenStudy (anonymous):

oh okay, let me see if I can figure it out from here

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Thanks a lot I figured it out! :)

OpenStudy (anonymous):

great!